• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 22:33:27 on October 4,1998

    Entry number 2353

    keyword=CODA problems fixed

    The ROC2 problem, which apparently was new and different from earlier
    problems, was fixed by cycling the power on the crate. It was hung up.

    By the way, the ELECTRON configuration would work if the ROC2 crate
    was turned off. The busy signals from ROC2 were stuck on, which prevented
    the ELECTRON config from taking data.